# Invisibility Cloak: Disappearance under Human Pose Estimation via Backdoor Attacks

### v1

@@ Abstract
Human Pose Estimation (HPE) has been widely applied in autonomous systems such as self-driving cars. However, the potential risks of HPE to adversarial attacks have not received comparable attention with image classification or segmentation tasks. Existing works on HPE robustness focus on misleading an HPE system to provide wrong predictions that still indicate some human poses. In this paper, we study the vulnerability of HPE systems to disappearance attacks, where the attacker aims to subtly alter the HPE training process via backdoor techniques so that any input image with some specific trigger will not be recognized as involving any human pose. As humans are typically at the center of HPE systems, such attacks can induce severe security hazards, e.g., pedestrians’ lives will be threatened if a self-driving car incorrectly understands the front scene due to disappearance attacks.
To achieve the adversarial goal of disappearance, we propose IntC, a general framework to craft Invisibility Cloak in the HPE domain. The core of our work lies in the design of target HPE labels that do not represent any human pose. In particular, we propose three specific backdoor attacks based on our IntC framework with different label designs. IntC-S and IntC-E, respectively designed for regression- and heatmap-based HPE techniques, concentrate the keypoints of triggered images in a tiny, imperceptible region. Further, to improve the attack’s stealthiness, IntC-L designs the target poisons to capture the label outputs of typical landscape images without a human involved, achieving disappearance and reducing detectability simultaneously. Extensive experiments demonstrate the effectiveness and generalizability of our IntC methods in achieving the disappearance goal. [^1] Our code is available at: https://github.com/TrustMLRG/IntC_HPE. By revealing the vulnerability of HPE to disappearance and backdoor attacks, we hope our work can raise awareness of the potential risks when HPE models are deployed in real-world applications.
1 Introduction
@@ 1.1 Our Work
- Through extensive experiments, our methods show effectiveness in achieving the disappearance goal for HPE models.
2 Related Works
Recall that our work aims to explore the vulnerabilities of HPE systems. Thus, we introduce HPE in this section, while other relevant related works are detailed in Appendix B due to space limit, including existing attacks against HPE, disappearance attacks, and backdoor attacks. As depicted in Figure 1, HPE predicts the 2D positions of a single person’s keypoints in an image, by which the human pose is determined. To be more specific, the input of an HPE model is an image $x$ containing a person and the output is the corresponding positions of the person’s keypoints $y$, which can be formulated as:
@@ 2 Related Works
Due to the advancement of deep learning, many works proposed to employ DNNs to improve the accuracy of HPE [57, 46, 18, 54]. In taxonomy, these studies can be categorized as either regression-based or heatmap-based methods. Specifically, regression-based methods directly predict the location of each keypoint. Toshev et al. first propose to use DNNs to solve this regression problem, i.e., DeepPose [57]. Inspired by DeepPose, some following works explore HPE as a regression task by taking advantage of the DNN development, e.g., adopting a re-parameterized and bone-based representation [55], using soft-argmax function to convert feature maps into keypoint locations [41] and involving transformers [31]. Another line of works proposed to transfer HPE into a heatmap-based problem [18, 62, 46, 12, 66, 54, 69, 70, 10], since heatmaps can preserve the spatial location information better and make the training process smoother [72]. In particular, each pixel value in a heatmap indicates the probability that the corresponding keypoint lies in the position. In addition, some recent works study the multi-person pose estimation [75, 26, 44, 17], i.e., human poses of multiple people are estimated in a single image. We conduct a primary study to explore the generalizability of our work in this more complex scenario in Section 6.4. Human poses in the 3D spatial location have also been explored in prior literature [2, 68, 43, 56]. Since this topic is less studied in the field and is regarded as the HPE technique followed by a matching module [7], we decided not to include it and left the exploration of our methods in 3D HPE as a future work.
@@
3 Threat Model
Adversarial Goal. Informally, the adversary aims to cause the person to “disappear” in the query image when predicted by the backdoored HPE model. As the first work that explores the disappearance in the HPE domain, we will conceptually introduce the notion of disappearance in this section and provide the specific objectives of our backdoor attacks in Section 4. In particular, disappearance is carried out not by erasing the pixels of the corresponding person in the input image but by misleading the target HPE model to output the label prediction that does not represent any person. By generalizing backdoor attacks to the HPE domain, we can explicitly connect triggers and the labels that indicate no person. Benefiting from the flexible HPE label space, our work proposes novel label designs to implement the disappearance task, which will be detailed in Section 4.
Adversarial Knowledge. Our work aims to accomplish the disappearance goal using backdoor techniques. Following existing backdoor works, we impose the following mild assumptions on the poisoning adversary, i.e., can access and manipulate a small portion of clean training dataset [52, 20]. To be more specific, the attacker first replaces a few clean samples for training the HPE model with the corresponding trigger-attached images and properly designed HPE labels. The model trained on such poisoned training dataset, denoted as backdoored HPE model throughout this paper, will provide a non-human prediction on any input image with the same trigger pattern. To maintain the reasonable level of attack stealthiness, the poisoning number and the attached trigger size are preferred to be as small as possible, which will be discussed in Section 7.1. Besides the poisoning access, we do not make any extra assumptions or requirements.
Although poisoning access is a commonly used setting in backdoor attacks, it would be helpful to discuss the practical feasibility of such an attack ability. The first potential possibility is intrusion techniques [34, 27]. Specifically, when an attacker evades or breaks the protection of the target model’s training data, they can easily modify the training data according to their expectations. Additionally, a large-scale dataset is normally collected from the internet [51, 35]. This decision provides the possibility of poisoning training data as the attacker can post their designed data online. Besides, cloud computing is a prevalent service where big companies provide individual users with computational resources, of which privacy and security concerns are worth heated discussions [61, 13]. For instance, a malicious service provider can easily access and modify the users’ stored data. Note that the above discussion serves to understand the access to poisoning training data. Besides, the exact skills will not be discussed in the paper, as our work focuses on the disappearance task in HPE. In the following, we will introduce how to achieve the disappearance goal with our novel label designs in Section 4.
4 Attack Methodology
4.1 Invisibility Cloak
@@
Problem Formulation. In the following, we will formulate the problem that our work aims to solve. As Figure 1 shows, IntC misleads the target model to think there is no person in an image by using the trigger. Specifically, given a clean training dataset $\mathcal{D}_{\text{tr}}\subseteq\mathcal{X}\times\mathcal{Y}$, where $\mathcal{X}$ denotes the input image space, and $\mathcal{Y}$ represents the label space of ground-truth keypoints. The attacker aims to poison a small part of $\mathcal{D}_{\text{tr}}$ with a poison rate $\alpha\in(0,1)$. Let $\mathcal{D}_{\text{c}}\subseteq\mathcal{D}_{\text{tr}}$ and $\mathcal{D}_{\text{p}}\subseteq\mathcal{D}_{\text{tr}}$ denote the set of clean examples that will be poisoned and the remaining unchanged set, respectively, where $|\mathcal{D}_{\text{p}}|=\alpha\cdot|\mathcal{D}_{\text{tr}}|$, $\mathcal{D}_{\text{c}}\cup\mathcal{D}_{\text{p}}=\mathcal{D}_{\text{tr}}$ and $\mathcal{D}_{\text{c}}\cap\mathcal{D}_{\text{p}}=\emptyset$. To generate the poisoned set $\mathcal{D}_{\text{p}}$, the attacker needs to employ a trigger injection function $\tau:\mathcal{X}\rightarrow\mathcal{X}$ and a label transformation function $\varphi:\mathcal{Y}\rightarrow\mathcal{Y}$. Specifically, the attacker first embeds a trigger, such as a small patch with specific patterns, to each input $(x_{\mathrm{p}},y_{\mathrm{p}})\in\mathcal{D}_{\text{p}}$ such that $x_{\mathrm{p}}^{\prime}=\tau(x_{\mathrm{p}})$ and then designs the corresponding label as $y_{\mathrm{p}}^{\prime}=\varphi(y_{\mathrm{p}})$ to indicate no person, where $y_{\mathrm{p}}$ denotes the original clean label of $x_{\mathrm{p}}$. [^2] This process is summarized in Algorithm 1 (Appendix D), followed by the details of IntC-B, IntC-S, IntC-E and IntC-L (Algorithm 2 – Algorithm 5). For ease of presentation, we denote the poisoned counterpart of $\mathcal{D}_{\text{p}}$ as:
@@
As mentioned in the challenge, simply selecting another label as the poison label cannot achieve the disappearance goal, since $y_{\mathrm{B}}$ still represents a person. In other words, even if the IntC-B attack can violate some HPE applications, e.g., abnormal monitoring, by selecting a benign human pose as $y_{\mathrm{B}}$, there is still a distance from disappearance as defined in Section 3. Consequently, we propose three novel label designs for solving the disappearance problem. Specifically, IntC-S and Int-E are separately designed for the regression- and heatmap-based HPE techniques, as described in Section 4.2 and Section 4.3. By setting the expected keypoints located in a tiny area, the HPE predictions on triggered images are hard to notice. Further, to improve the attack stealthiness, we propose IntC-L by capturing the information of landscape images as the desired label, which will be introduced in Section 4.4. In that case, triggered images will be regarded as landscape images, i.e., no person is involved.
4.2 IntC Against Regression-Based HPE
@@
4.3 IntC Against Heatmap-Based HPE
@@
4.4 Improvement on Attack Steathiness
@@
Figure 3: Distributions of different input images by projecting predicted HPE labels to 2-dimensional space via t-SNE.
@@
4.5 Summary of Our Attacks
Table 1 summarizes the settings of the proposed baseline and our IntC attacks, including whether satisfying the disappearance requirement, application scopes, label designs, and required extra knowledge. We can see that except for the baseline IntC-B, our IntC attacks (i.e., IntC-S, IntC-E and IntC-L) all support the disappearance task. Further, as only the label design of IntC-E is specific to heatmap-based HPE models, the others are all applicable to both the regression- and heatmap-based HPE techniques. In addition, our IntC-S and IntC-E attacks do not require extra knowledge. Meanwhile, IntC-B requires a new clean data point $(x_{\mathrm{B}},y_{\mathrm{B}})$, and IntC-L requires the black-box access to a clean pre-trained HPE model $M_{\theta_{\mathrm{c}}}$ and a set of landscape images $\mathcal{S}_{\mathrm{L}}$.
Table 1: The settings of the baseline IntC-B and our IntC attacks, i.e., IntC-S, IntC-E and IntC-L, where “Regression” and “Heatmap” respectively represent the regression- and heatmap-based HPE techniques, and $\mathcal{B}(\cdot)$ represent the black-box access.

### v2

@@ Abstract
Despite being significant in autonomous systems, Human Pose Estimation (HPE)’s potential risks to adversarial attacks have not received comparable attention with image classification or segmentation tasks. In this paper, we study the vulnerability of HPE systems to disappearance attacks, where the attacker aims to subtly alter the HPE training process via backdoor techniques so that any input image with some specific trigger will not be recognized as involving any human pose. As humans are typically at the center of HPE systems, a successful attack will severely threaten pedestrians’ lives if a self-driving car incorrectly understands the front scene. To achieve the adversarial goal of disappearance, we propose IntC, a general framework to craft an invisibility cloak in the HPE domain. By designing target HPE labels that do not represent any human pose, we propose three specific backdoor attacks based on our IntC framework. IntC-S and IntC-E, respectively designed for regression- and heatmap-based HPE techniques, concentrate the keypoints of triggered images in a tiny, imperceptible region. Further, to improve the attack’s stealthiness, IntC-L designs the target poisons to capture the label outputs of typical landscape images without a human involved, achieving disappearance and reducing detectability simultaneously. Extensive experiments demonstrate the effectiveness and generalizability of our IntC methods in achieving the disappearance goal. By revealing the vulnerability of HPE to disappearance and backdoor attacks, we hope our work can raise awareness of the potential risks when HPE models are deployed in real-world applications.
I Introduction
@@ I-A Our Work
  • Through extensive experiments, our methods show effectiveness and generalizability in achieving our disappearance goal for HPE models.
II Background and Related Works
Human Pose Estimation. As depicted in Figure 1a, HPE predicts the 2D positions of a single person’s keypoints in an image, by which the human pose is determined. To be more specific, the input of an HPE model is an image $x$ containing a person, and the output is the corresponding positions of the person’s keypoints $y$, which can be formulated as:
@@ II Background and Related Works
Due to the advancement of deep learning, many works have proposed to employ DNNs to improve the accuracy of HPE [1, 29, 17, 18]. In taxonomy, these studies can be categorized as either regression-based or heatmap-based methods. Specifically, regression-based methods directly predict the location of each keypoint. Toshev et al. first propose to use DNNs to solve this regression problem, i.e., DeepPose [1]. Inspired by DeepPose, some following works explore HPE as a regression task by taking advantage of the DNN development, e.g., adopting a re-parameterized and bone-based representation [30], using soft-argmax function to convert feature maps into keypoint locations [31], and involving transformers [32]. Another line of work proposed to transfer HPE into a heatmap-based problem [17, 33, 29, 34, 35, 18, 36, 37, 38], as heatmaps can preserve the spatial location information better and make the training process smoother [39]. In particular, each pixel value in a heatmap indicates the probability that the corresponding keypoint lies in the position. In addition, some recent works study the multi-person pose estimation [40, 41, 42, 19], i.e., human poses of multiple people are estimated in a single image. A primary study is conducted to explore the generalizability of our work in this more complex scenario in Section VI-D. Human poses in the 3D spatial location have also been explored in prior literature [43, 44, 45, 46]. As this topic is less studied in the field and is regarded as the HPE technique followed by a matching module [47], we decided not to include it and left the exploration in 3D HPE as future work.
Disappearance Attacks. The concepts of disappearance attacks and invisibility cloaks have been investigated in various computer-vision tasks [48, 49, 50, 51], including object detection and facial recognition. Existing works manage to achieve the disappearance goal by misclassification solutions. For instance, a target model is misled to predict an incorrect category (i.e., identity) for the detected object (i.e., a human face image), where the predicted label is expected to be different from the ground truth. However, a different predicted HPE label still suggests that a person is in the captured image, which does not support our disappearance goal. This gap is derived from the fact that HPE is not a simple classification task, which motivates our work that aims to design proper HPE labels that do not represent any human pose.
Backdoor Attacks. Compared to evasion attacks, backdoor attacks have the advantage of constructing an adversarial connection between a certain trigger and any desired label by poisoning strategies while minimizing the influence on the predictions of clean inputs. The majority of existing backdoor attacks focused on classification tasks [14, 52, 53, 15, 54, 55, 56, 57], where an adversary can strategically attach triggers to a few training data and modify the corresponding labels. When trained on the poisoned data, the target model can be misled to predict the expected class on any triggered image. In contrast, our work generalizes the backdoor technique to the HPE domain, aiming to achieve the disappearance goal. The disappearance feasibility is derived from the flexible HPE label space. Specifically, each person has several keypoints, each of which can be located at any position in an image. In that case, by probably designing poison labels, a backdoored HPE model can provide a prediction that indicates no person in a triggered image. In other words, the core of the backdoor attack against HPE is how to design the desired labels. By proposing novel label designs, our work brings a new attack surface to explore the security risks in the HPE domain.
III Threat Model
Adversarial Goal. Informally, the adversary aims to cause the person to “disappear” in the query image when predicted by the backdoored HPE model. As the first work that explores the disappearance in the HPE domain, we will conceptually introduce the notion of disappearance in this section and provide the specific objectives of our backdoor attacks in Section IV. In particular, disappearance is carried out not by erasing the pixels of the corresponding person in the input image but by misleading the target HPE model to output the label prediction that does not represent any person. By generalizing backdoor attacks to the HPE domain, we can explicitly connect triggers and the labels that indicate no person. Benefiting from the flexible HPE label space, our work proposes novel label designs to implement the disappearance task, which will be detailed in Section IV.
@@ III Threat Model
Adversarial Knowledge. Our work aims to accomplish the disappearance goal using backdoor techniques. Following existing backdoor works, we impose the following mild assumptions on the poisoning adversary, i.e., can access and manipulate a small portion of clean training dataset [59, 60]. To be more specific, the attacker first replaces a few clean samples for training the HPE model with the corresponding trigger-attached images and properly designed HPE labels. The model trained on such poisoned training dataset, denoted as backdoored HPE model throughout this paper, will provide a non-human prediction on any input image with the same trigger pattern. To maintain the reasonable level of attack stealthiness, the poisoning number and the attached trigger size are preferred to be as small as possible, which will be discussed in Section VII-A. Besides the poisoning access, we do not make any extra assumptions or requirements.
Although poisoning access is a commonly used setting in backdoor attacks, it would be helpful to discuss the practical feasibility of such an attack ability. The first potential possibility is intrusion techniques [61, 62]. Specifically, when an attacker evades or breaks the protection of the target model’s training data, they can easily modify the training data according to their expectations. Additionally, a large-scale dataset is normally collected from the internet [63, 20]. This decision provides the possibility of poisoning training data as the attacker can post their designed data online. Besides, cloud computing is a prevalent service where big companies provide individual users with computational resources, of which privacy and security concerns are worth heated discussions [64, 65]. For instance, a malicious service provider can easily access and modify the users’ stored data. Note that the above discussion serves to understand the access to poisoning training data. Besides, the exact skills will not be discussed in the paper, as our work focuses on the disappearance task in HPE. In the following, we will introduce how to achieve the disappearance goal with our novel label designs in Section IV.
IV Attack Methodology
IV-A Invisibility Cloak
@@
Algorithm 1 Training Data Poisoning of Invisibility Cloak
  $\mathcal{D}_{\text{c}}$, $\mathcal{D}_{\text{p}}$ $\leftarrow$ Split training data $\mathcal{D}_{\text{tr}}$ by poison rate $\alpha$
  if “IntC-B” then
   $\varphi(\cdot)$ $\leftarrow$ $\text{IntC-B}(\cdot)$ $\triangleright$ Algorithm 2.
  else if “IntC-S” then
   $\varphi(\cdot)$ $\leftarrow$ $\text{IntC-S}(\cdot)$ $\triangleright$ Algorithm 3.
  else if “IntC-E” then
   $\varphi(\cdot)$ $\leftarrow$ $\text{IntC-E}(\cdot)$ $\triangleright$ Algorithm 4.
  else if “IntC-L” then
   $\varphi(\cdot)$ $\leftarrow$ $\text{IntC-L}(\cdot)$ $\triangleright$ Algorithm 5.
  $\mathcal{X}^{\prime}_{\mathrm{p}}$ $\leftarrow$ $\tau(\mathcal{X}_{\mathrm{p}})$ $\triangleright$ Inject triggers by $\tau(\cdot)$.
  $\mathcal{Y}^{\prime}_{\mathrm{p}}$ $\leftarrow$ $\varphi(\mathcal{Y}_{\mathrm{p}})$ $\triangleright$ Transform labels by $\varphi(\cdot)$.
  $\mathcal{D}^{\prime}_{\text{p}}$ $\leftarrow$ $\mathcal{X}^{\prime}_{\mathrm{p}}\times\mathcal{Y}^{\prime}_{\mathrm{p}}$
Problem Formulation. In the following, we will formulate the problem that our work aims to solve. As Figure 1b shows, IntC misleads the target model to think there is no person in an image by using the trigger. Specifically, given a clean training dataset $\mathcal{D}_{\text{tr}}\subseteq\mathcal{X}\times\mathcal{Y}$, where $\mathcal{X}$ denotes the input image space, and $\mathcal{Y}$ represents the label space of ground-truth keypoints. The attacker aims to poison a small part of $\mathcal{D}_{\text{tr}}$ with a poison rate $\alpha\in(0,1)$. Let $\mathcal{D}_{\text{c}}\subseteq\mathcal{D}_{\text{tr}}$ and $\mathcal{D}_{\text{p}}\subseteq\mathcal{D}_{\text{tr}}$ denote the set of clean examples that will be poisoned and the remaining unchanged set, respectively, where $|\mathcal{D}_{\text{p}}|=\alpha\cdot|\mathcal{D}_{\text{tr}}|$, $\mathcal{D}_{\text{c}}\cup\mathcal{D}_{\text{p}}=\mathcal{D}_{\text{tr}}$ and $\mathcal{D}_{\text{c}}\cap\mathcal{D}_{\text{p}}=\emptyset$. To generate the poisoned set $\mathcal{D}_{\text{p}}$, the attacker needs to employ a trigger injection function $\tau:\mathcal{X}\rightarrow\mathcal{X}$ and a label transformation function $\varphi:\mathcal{Y}\rightarrow\mathcal{Y}$. Specifically, the attacker first embeds a trigger, such as a small patch with specific patterns, to each input $(x_{\mathrm{p}},y_{\mathrm{p}})\in\mathcal{D}_{\text{p}}$ such that $x_{\mathrm{p}}^{\prime}=\tau(x_{\mathrm{p}})$ and then designs the corresponding label as $y_{\mathrm{p}}^{\prime}=\varphi(y_{\mathrm{p}})$ to indicate no person, where $y_{\mathrm{p}}$ denotes the original clean label of $x_{\mathrm{p}}$. The process is summarized in Algorithm 1, followed by the details of IntC-B, IntC-S, IntC-E, and IntC-L (Algorithm 2 – Algorithm 5). For ease of presentation, we denote the poisoned counterpart of $\mathcal{D}_{\text{p}}$ as:
@@
Algorithm 2 The algorithm of IntC-B
function IntC-B(To-be-poisoned labels $\mathcal{Y}_{\mathrm{p}}$)
  $y_{\mathrm{B}}$ $\leftarrow$ A new clean data point $(x_{\mathrm{B}},y_{\mathrm{B}})$ $\triangleright$ $x_{\mathrm{B}}\notin\mathcal{D}_{\text{p}}$.
  for $i\leftarrow 1$ to $|\mathcal{Y}_{\mathrm{p}}|$ do
   $y_{\mathrm{p}}^{i}$ $\leftarrow$ $y_{\mathrm{B}}$ $\triangleright$ $\mathcal{Y}_{\mathrm{p}}=\{y_{\mathrm{p}}^{i}\}_{i=1,\cdots,|\mathcal{Y}_{\mathrm{p}}|}$.
  return Poisoned labels $\mathcal{Y}^{\prime}_{\mathrm{p}}$
@@
As mentioned in the challenge, simply selecting another label as the poison label cannot achieve the disappearance goal, since $y_{\mathrm{B}}$ still represents a person. In other words, even if the IntC-B attack can violate some HPE applications, e.g., abnormal monitoring, by selecting a benign human pose as $y_{\mathrm{B}}$, there is still a distance from disappearance as defined in Section III. Consequently, we propose three novel label designs for solving the disappearance problem. Specifically, IntC-S and Int-E are separately designed for the regression- and heatmap-based HPE techniques, as described in Section IV-B and Section IV-C. By setting the expected keypoints located in a tiny area, the HPE predictions on triggered images are hard to notice. Further, to improve the attack stealthiness, we propose IntC-L by capturing the information of landscape images as the desired label, which will be introduced in Section IV-D. In that case, triggered images will be regarded as landscape images, i.e., no person is involved.
IV-B IntC Against Regression-Based HPE
@@
Algorithm 3 The algorithm of IntC-S
function IntC-S(To-be-poisoned labels $\mathcal{Y}_{\mathrm{p}}$)
  $\{(a,b)\times n\}$ $\leftarrow$ Decide a desired location $(a,b)$
  $\triangleright$ $n$ represents the joint number of a person.
  for $i\leftarrow 1$ to $|\mathcal{Y}_{\mathrm{p}}|$ do
   $y_{\mathrm{p}}^{i}$ $\leftarrow$ $\{(a,b)\times n\}$ $\triangleright$ $\mathcal{Y}_{\mathrm{p}}=\{y_{\mathrm{p}}^{i}\}_{i=1,\cdots,|\mathcal{Y}_{\mathrm{p}}|}$.
  return Poisoned labels $\mathcal{Y}^{\prime}_{\mathrm{p}}$
@@
IV-C IntC Against Heatmap-Based HPE
@@
Algorithm 4 The algorithm of IntC-E
function IntC-E(To-be-poisoned labels $\mathcal{Y}_{\mathrm{p}}$)
  $\mathbf{0}$ $\leftarrow$ Set $n$ empty heatmaps
  for $i\leftarrow 1$ to $|\mathcal{Y}_{\mathrm{p}}|$ do
   $y_{\mathrm{p}}^{i}$ $\leftarrow$ $\mathbf{0}$ $\triangleright$ $\mathcal{Y}_{\mathrm{p}}=\{y_{\mathrm{p}}^{i}\}_{i=1,\cdots,|\mathcal{Y}_{\mathrm{p}}|}$.
  return Poisoned labels $\mathcal{Y}^{\prime}_{\mathrm{p}}$
@@
IV-D Improvement on Attack Steathiness
@@
Algorithm 5 The algorithm of IntC-L
function IntC-L(To-be-poisoned labels $\mathcal{Y}_{\mathrm{p}}$)
  $\mathcal{Y}_{\mathrm{L}}$ $\leftarrow$ Predictions on landscape images $\mathcal{M}_{\theta_{\mathrm{c}}}(\mathcal{X}_{\mathrm{L}})$
  $\triangleright$ $\mathcal{M}_{\theta_{\mathrm{c}}}$ is a clean HPE model.
  $\bar{y}_{\mathrm{L}}$ $\leftarrow$ $\frac{1}{|\mathcal{Y}_{\mathrm{L}}|}\sum_{i=1}^{\mathcal{Y}_{\mathrm{L}}}y_{\mathrm{L}}^{i}$ $\triangleright$ $\mathcal{Y}_{\mathrm{L}}=\{y_{\mathrm{L}}^{i}\}_{i=1,\cdots,|\mathcal{Y}_{\mathrm{L}}|}$.
  for $i\leftarrow 1$ to $|\mathcal{Y}_{\mathrm{p}}|$ do
   $y_{\mathrm{p}}^{i}$ $\leftarrow$ $\bar{y}_{\mathrm{L}}$ $\triangleright$ $\mathcal{Y}_{\mathrm{p}}=\{y_{\mathrm{p}}^{i}\}_{i=1,\cdots,|\mathcal{Y}_{\mathrm{p}}|}$.
  return Poisoned labels $\mathcal{Y}^{\prime}_{\mathrm{p}}$
@@
Figure 3: Distributions of different input images by projecting predicted HPE labels to a 2-dimensional space via t-SNE.
@@
IV-E Summary of Our Attacks
Table I summarizes the settings of the proposed baseline and our IntC attacks, including whether satisfying the disappearance requirement, application scopes, label designs, and required extra knowledge. We can see that except for the baseline IntC-B, our IntC attacks (i.e., IntC-S, IntC-E and IntC-L) all support the disappearance task. Further, as only the label design of IntC-E is specific to heatmap-based HPE models, the others are all applicable to both the regression- and heatmap-based HPE techniques. In addition, our IntC-S and IntC-E attacks do not require extra knowledge. Meanwhile, IntC-B requires a new clean data point $(x_{\mathrm{B}},y_{\mathrm{B}})$, and IntC-L requires the black-box access to a clean pre-trained HPE model $M_{\theta_{\mathrm{c}}}$ and a set of landscape images $\mathcal{S}_{\mathrm{L}}$.
TABLE I: The settings of the baseline IntC-B and our IntC attacks, i.e., IntC-S, IntC-E and IntC-L, where “Regression” and “Heatmap” respectively represent the regression- and heatmap-based HPE techniques, and $\mathcal{B}(\cdot)$ represent the black-box access.